# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

### v2

@@ Appendix F Topographic ICA (TICA)
F.1 Overview of TICA
$\sigma_{\ell}$ $=\phi\left(\sum_{m=1}^{d}h_{\ell m}u_{m}\right),$ (25)
$s_{\ell}$ $=\sigma_{\ell}z_{\ell},$ (26)
where $\phi$ is some nonlinear function, $h_{\ell m}$ is a symmetric neighborhood relation, and $z_{\ell}$ are mutually independent random variables. To estimate the decomposition matrix $\mathbf{W}=[\mathbf{w}_{1},\cdots,\mathbf{w}_{d}]^{\top}\in\mathbb{R}^{d\times d}$, where $\mathbf{w}_{\ell}\in\mathbb{R}^{d}$, such that $\mathbf{s}=\mathbf{W}\mathbf{x}$ from the mixed signals $\mathbf{x}\in\mathbb{R}^{d}$, TICA maximize the log likelihood
The optimization of (28) can be done by gradient descent:
Note that we have to apply orthonormalization
$\mathbf{W}\leftarrow(\mathbf{WW}^{\top})^{-1/2}\mathbf{W}$ (30)
$\mathbf{S}=\mathbf{X}\mathbf{W}^{\top}.$ (31)
F.2 Experiments
As in classic ICA, we transform the $300$-dimensional GloVe by TICA for our experiments. Following the setup of the 1-D experiment in Hyvärinen et al. (2001a), a neighborhood relation was defined by convolving a rectangular-shaped filter $(\ldots,0,0,0,\overbrace{1,1,\ldots,1,1}^{\mathsf{width}~{}\text{times}},0,0,0,\ldots)$ twice with itself. Hyvärinen et al. (2001a) defined the hyperparameter $\mathsf{width}=5$ for 20 components, so for 300 components, we consider two cases where the width is scaled logarithmically or linearly to $\mathsf{width}=9$ and $\mathsf{width}=75$. These are denoted as TICA9 and TICA75, respectively. We performed $10{,}000$ iterations of gradient descent in both settings.
F.2.1 Qualitative observation
@@ F.2.2 Scatterplots of Table 10
Figs. 17 and 18 show the scatterplots of the two-dimensional projections for the the axes of the TICA9-transformed embeddings and the TICA75-transformed embeddings in Tables 10(a) and 10(b). In these examples, for both TICA9 and TICA75, the interpretability of the embedding space from the scatterplots decreases as the top words of the axes show significant movement or approach the origin. This property is particularly noticeable in TICA75. For example, in Fig. 18, the top words of the 159th axis are far from the axis: surged and soared are concentrated near the 156th axis, and leapt and leaped are concentrated near the origin.
@@
Table 11: The values of $d_{I}$ for Figs. 1718 and 19, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
@@
Table 12: The values of $c_{I}$ for Figs. 1718 and 19, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
@@ F.3.1 Scatterplots
This section compares the scatterplots of the TICA9-transformed embeddings in Figure 17 and the TICA75-transformed embeddings in Figure 18 with those of Axis Tour. Since TICA9, TICA75, and Axis Tour yield different embeddings, we also compare them based on Skewness Sort.
First, similar to Figs. 17 and 18, we use 300-dimensional GloVe and show the scatterplots for comparison232323 Note that Figs. 1 and  5 show the scatterplots for the illustrative examples, so it is unfair to compare them directly with those of TICA. in Fig. 19, which shows the two-dimensional projections for the axes of the Axis Tour embeddings.
Then Table 11 shows the values of $d_{I}$ defined in Appendix A.3 for Figs. 1718, and 19, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. In these examples, Axis Tour shows larger $d_{I}$ values and larger differences with Skewness Sort compared to TICA9 and TICA75. These results are also observed for the average values. In addition, we can see the cases where $d_{I}$ is smaller than that of Skewness Sort in both TICA9 and TICA75, such as Figures 17 and 18.
Table 12 shows the values of $c_{I}$ defined in Appendix A.3 for Figs. 1718, and 19, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. In these examples, TICA9 shows the $c_{I}$ values comparable to Axis Tour. In fact, their average values are almost the same. However, TICA9 shows larger $c_{I}$ values even when Skewness Sort is performed on $I$. This suggests that for the TICA-transformed embeddings, the axes with similar meanings tend to cluster locally rather than being ordered so that the meanings are continuous. This hypothesis could explain why, in the scatterplots such as Figures 17 and 18, the top words of the axes show significant movement or approach the origin.
@@ F.3.2 Cosine similarity and higher-order correlation
Fig. 20 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. Similar to Axis Tour, The distribution for TICA shifts towards a more positive mean than that of Random Order. Note that while TICA9 has an average close to that of Axis Tour, it has a significantly higher variance of cosine similarity.
To highlight this, Figure 21 shows scatterplots of cosine similarities between adjacent axis embeddings. The variance is $0.05$ for Random Order, $0.12$ for Axis Tour, $0.25$ for TICA9, and 0.15 for TICA75. The variance for TICA9 is significantly larger than that for Axis Tour, which indicates that the semantic continuity of the axes changes more drastically in the TICA9-transformed embeddings.
Fig. 23 shows the histograms of higher-order correlations between adjacent axes of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. We can see that the average higher-order correlation for TICA is higher than that for Axis Tour, reflecting the learning settings of TICA.
@@ F.3.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Fig. 22 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the Axis Tour embeddings, and TICA-transformed embeddings. Axis Tour outperformed both TICA and TICA Projection for most dimensions in each task. Note that similar to the results for Random Order and Skewness Sort in Appendix E.4, TICA Projection shows performance improvements over TICA in most dimensions. These results demonstrate the utility of projection-based dimensionality reduction even in TICA, an algorithm that relaxes the assumption of statistical indepence for ICA.

### v3

@@ F.2.2 Scatterplots of Table 10
Figures 18 and 19 show the scatterplots of the two-dimensional projections for the the axes of the TICA9-transformed embeddings and the TICA75-transformed embeddings in Tables 10(a) and 10(b). In these examples, for both TICA9 and TICA75, the interpretability of the embedding space from the scatterplots decreases as the top words of the axes show significant movement or approach the origin. This property is particularly noticeable in TICA75. For example, in Fig. 19, the top words of the 159th axis are far from the axis: surged and soared are concentrated near the 156th axis, and leapt and leaped are concentrated near the origin.
@@
Table 11: The values of $d_{I}$ for Figs. 1819 and 20, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
@@
Table 12: The values of $c_{I}$ for Figs. 1819 and 20, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
@@ F.3.1 Scatterplots
This section compares the scatterplots of the TICA9-transformed embeddings in Fig. 18 and the TICA75-transformed embeddings in Fig. 19 with those of Axis Tour. Since TICA9, TICA75, and Axis Tour yield different embeddings, we also compare them based on Skewness Sort.
First, similar to Figs. 18 and 19, we use 300-dimensional GloVe and show the scatterplots for comparison232323 Note that Figs. 1 and  6 show the scatterplots for the illustrative examples, so it is unfair to compare them directly with those of TICA. in Fig. 20, which shows the two-dimensional projections for the axes of the Axis Tour embeddings.
Then Table 11 shows the values of $d_{I}$ defined in Appendix A.3 for Figs. 1819, and 20, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. In these examples, Axis Tour shows larger $d_{I}$ values and larger differences with Skewness Sort compared to TICA9 and TICA75. These results are also observed for the average values. In addition, we can see the cases where $d_{I}$ is smaller than that of Skewness Sort in both TICA9 and TICA75, such as Figs. 18 and 19.
Table 12 shows the values of $c_{I}$ defined242424 For the TICA9- and TICA75-transformed embeddings, as with Axis Tour embeddings, $c_{I}$ can be computed by the cosine similarities of adjacent axis embeddings. in Appendix A.3 for Figs. 1819, and 20, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. In these examples, TICA9 shows the $c_{I}$ values comparable to Axis Tour. In fact, their average values are almost the same. However, TICA9 shows larger $c_{I}$ values even when Skewness Sort is performed on $I$. This suggests that for the TICA-transformed embeddings, the axes with similar meanings tend to cluster locally rather than being ordered so that the meanings are continuous. This hypothesis could explain why, in the scatterplots such as Figs. 18 and 19, the top words of the axes show significant movement or approach the origin.
@@ F.3.2 Cosine similarity and higher-order correlation
Figure 21 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. Similar to Axis Tour, The distribution for TICA shifts towards a more positive mean than that of Random Order. Note that while TICA9 has an average close to that of Axis Tour, it has a significantly higher variance of cosine similarity.
To highlight this, Fig. 22 shows scatterplots of cosine similarities between adjacent axis embeddings. The variance is $0.05$ for Random Order, $0.12$ for Axis Tour, $0.25$ for TICA9, and 0.15 for TICA75. The variance for TICA9 is significantly larger than that for Axis Tour, which indicates that the semantic continuity of the axes changes more drastically in the TICA9-transformed embeddings.
Figure 24 shows the histograms of higher-order correlations between adjacent axes of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. We can see that the average higher-order correlation for TICA is higher than that for Axis Tour, reflecting the learning settings of TICA.
@@ F.3.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Role Prompt
system You are an excellent NLP annotator. Your response should be in JSON format with the key ‘choice’.
user Which of the following words are related to the words [top1 word,$\ldots$, top10 word]. Answer A or B.
A. [Axis Tour top1 word,$\ldots$, Axis Tour top10 word]
B. [Skewness Sort top1 word,$\ldots$, Skewness Sort top10 word]
Table 13: Prompts for the GPT models. Since the Axis Tour embeddings and the Skewness Sort embeddings differ solely in the order of the axes, they share $d\,(=300)$ common axes. The top 10 words for each common axis are denoted as [top1 word,$\ldots$, top10 word]. We focus on the next axis of Axis Tour and the next axis of Skewness Sort, based on the common axis. The top 10 words for the Axis Tour axis are denoted as [Axis Tour top1 word,$\ldots$, Axis Tour top10 word], while the top 10 words for the Skewness Sort axis are denoted as [Skewness Sort top1 word,$\ldots$, Skewness Sort top10 word]. We then use the top words for the prompt.
Model Version
GPT-3.5 Turbo gpt-3.5-turbo-0125
GPT-4 Turbo gpt-4-turbo-2024-04-09
GPT-4o gpt-4o-2024-05-13
GPT-4o mini gpt-4o-mini-2024-07-18
Table 14: Version of each GPT model.
Figure 23 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the Axis Tour embeddings, and TICA-transformed embeddings. Axis Tour outperformed both TICA and TICA Projection for most dimensions in each task. Note that similar to the results for Random Order and Skewness Sort in Appendix E.4, TICA Projection shows performance improvements over TICA in most dimensions. These results demonstrate the utility of projection-based dimensionality reduction even in TICA, an algorithm that relaxes the assumption of statistical indepence for ICA.
Appendix G Details of the quantitative evaluation of semantic continuity by GPT models
This section provides details on the GPT models, accessed via the OpenAI API, and the prompts used in Section 5.2.2.
The versions of the GPT models used in the experiments are listed in Table 14. The prompts used for the experiments are shown in Table 13. In Fig. 3, the responses obtained from each model using these prompts are aggregated and shown. Note that GPT-4 Turbo failed to provide responses relevant to either Axis Tour or Skewness Sort for two queries.